Given f(x)=-2x^2 + x-5, find f(x+3).
Answer by jim_thompson5910(35256)

f(x) = -2x^2 + x-5


f(x) = -2x^2 + 1x-5


f(x+3) = -2(x+3)^2 + 1(x+3)-5


f(x+3) = -2(x^2+6x+9) + 1(x+3)-5


f(x+3) = -2(x^2) - 2(6x) - 2(9) + 1(x) + 1(3) - 5


f(x+3) = -2x^2 - 12x - 18 + x + 3 - 5


f(x+3) = -2x^2 - 11x - 20


If you need more help, feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim